# Efficient Dataset Distillation via Diffusion-Driven Patch Selection for Improved Generalization

### v3

@@ Abstract
Dataset distillation offers an efficient way to reduce memory and computational costs by optimizing a smaller dataset with performance comparable to the full-scale original. However, for large datasets and complex deep networks (e.g., ImageNet-1K with ResNet-101), the extensive optimization space limits performance, reducing its practicality. Recent approaches employ pre-trained diffusion models to generate informative images directly, avoiding pixel-level optimization and achieving notable results. However, these methods often face challenges due to distribution shifts between pre-trained models and target datasets, along with the need for multiple distillation steps across varying settings. To address these issues, we propose a novel framework orthogonal to existing diffusion-based distillation methods, leveraging diffusion models for selection rather than generation. Our method starts by predicting noise generated by the diffusion model based on input images and text prompts (with or without label text), then calculates the corresponding loss for each pair. With the loss differences, we identify distinctive regions of the original images. Additionally, we perform intra-class clustering and ranking on selected patches to maintain diversity constraints. This streamlined framework enables a single-step distillation process, and extensive experiments demonstrate that our approach outperforms state-of-the-art methods across various metrics.
@@ 1 Introduction
Figure 1: Comparison of various methods for dataset distillation using diffusion models. (a) Fine-tuning a diffusion model pre-trained on ImageNet-1K to directly generate images. (b) Calculating the clustering center of the original images in latent space to generate images. (c) Identifying the regions in the real images most strongly associated with label text based on the implicit data distribution learned by the diffusion model. The images generated by our method more effectively represent the feature distribution of the original dataset.
@@
The rapid advancement of deep learning has driven impressive performance gains through increasingly deeper and more complex models trained on large-scale datasets [9, 5]. However, training these models demands significant memory and computational resources. Dataset distillation has recently emerged as an effective approach to data compression [17], offering substantial reductions in resource consumption. By optimizing images in pixel space through information matching (e.g., gradient matching), these optimization-based methods produce smaller synthetic datasets that can achieve comparable performance to the original data [37, 36, 1]. Despite these benefits, the vast pixel-level optimization space restricts the feasibility of these methods to small-scale datasets, e.g., CIFAR-10/100 [15], thus limits their generalization.
A recent approach, SRe2[35], scales optimization-based methods to larger datasets (e.g., ImageNet-1K [4]) by using a pre-trained classifier to guide dataset synthesis. However, its strong dependence on proxy models during optimization often hinders the distilled dataset’s generalization, limiting broader applicability. Another critical limitation arises when handling various distillation configurations, such as different subsets of large datasets. Previous optimization-based methods typically incur substantial redundant costs due to repeated distillation requirements. Although recent works [11, 10] aim to address this issue, their applicability remains confined to small-scale datasets.
@@
To address the limitations of current dataset distillation methods, recent approaches [31, 7, 30] that avoid direct image optimization have shown promise for scaling dataset distillation to larger datasets and more complex model architectures. RDED [31] leverages a pre-trained classifier (e.g., ResNet-18 [9]) to select patches randomly cropped from real images and then concatenate them into synthetic images. While RDED achieves substantial performance improvements, it requires careful model selection and retraining of a classifier tailored to the specific dataset. Moreover, concatenated images often lack complete class-relevant features, especially for complex models with higher learning capacity. Minimax [7] addresses distillation by pruning the DiT model [22] to generate synthetic datasets using regularization based on dataset representativeness and diversity. However, Minimax relies on a diffusion model pre-trained on ImageNet-1k [4], requiring fine-tuning for different target datasets, which limits flexibility. D4[30] applies a pre-trained text-to-image Latent Diffusion Model (LDM)[23] to generate synthetic datasets by clustering original images in latent space. However, without constraints on distributional differences between the LDM’s pre-training data (e.g., LAION[26]) and the target dataset (e.g., ImageNet-1K), D4M can produce class-irrelevant or noisy images.
@@
To overcome these limitations, we propose a novel dataset distillation paradigm that leverages diffusion model’s prior to select the most informative image patches. Specifically, for each input image from the original dataset, we first predict the noise map using the pre-trained LDM, with and without the label text prompt, calculating losses for each scenario. The differential loss quantifies representativeness, enabling us to crop patches most relevant to the label guided by the target dataset’s data distribution. Additionally, we can effectively constrain the implicit data distribution learned by the diffusion model by using the pre-trained LDM to identify class-pertinent regions.
@@
To mitigate the inherent diversity of diffusion models, we apply clustering to capture visual features most representative of each class. Specifically, we compute diffusion features for all patches inspired by recent work [32], allowing for efficient clustering and ranking of cropped patches. Fig. 1 illustrates a comparison between our method and previous diffusion-based generative distillation methods. Extensive experiments demonstrate that our approach achieves state-of-the-art (SOTA) performance on large-scale datasets, supporting an efficient, unrestricted one-step distillation process.
@@
  • Our proposed paradigm enables a highly efficient, one-step distillation process, eliminating the need for model training and fine-tuning. Additionally, it avoids repetitive distillation across subclass combinations or compression ratios, which substantially broadens the potential applications and flexibility of dataset distillation.
  • Extensive experiments demonstrate that our method consistently surpasses existing training-free approaches, particularly on large-scale datasets, underscoring the effectiveness and robustness of our framework in various challenging settings.
@@ 2.1 Dataset Distillation
Dataset distillation was initially regarded as a meta-learning problem [34]. It involves minimizing the loss function of the synthetic dataset through a model trained on this dataset. Since then, several approaches have been proposed to enhance the performance of dataset distillation. One category of methods utilizes ridge regression models to approximate the inner loop optimization [20, 21, 41]. To address the issue of data distribution bias in meta-learning, another category of methods selects an information space and computes proxy information to align the synthetic dataset with the original dataset. DC [39] and DSA [37] match the weight gradients of models trained on the real and synthetic dataset, respectively, while DM [38], CAFE [33] and DataDAM [25] focus on using feature distance between the two datasets as matching targets. MTT [1] and TESLA [3] match the model parameter trajectories obtained from training on the real dataset and the synthetic dataset.
Recent methods propose the decoupled distillation strategy and successfully extend the dataset distillation application to large-scale datasets. SRe2[35] leverages a pre-trained classifier to compute the cross-entropy loss of the synthetic dataset for updates, while RDED [31] calculates the confidence of randomly cropped patches from the original images with a pre-trained classifier and concatenates them together to form the synthetic dataset. Recent methods [2, 40] have introduced GAN as a parameterization technique. While with the development of diffusion models, Minimax [7] and D4[30] introduce the diffusion models to directly generate synthetic datasets, achieving promising performance. However, existing diffusion-based methods are constrained by the implicit data distribution learned by diffusion models, leading to significant distributional bias in the synthetic dataset. In addition, aforementioned methods still fail to complete the distillation process in one pass for all settings. To address these issues, we utilize diffusion model to actively identify the most class-relevant regions in the original images, enabling an efficient dataset distillation process.
@@ 2.2 Diffusion Model
As a class of generative approaches, diffusion-based frameworks learn to progressively convert Gaussian-distributed noise vectors $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into samples matching target distributions $\mathcal{X}$, demonstrating remarkable effectiveness across diverse applications [42, 43, 28]. These models employ a neural network $\epsilon_{\theta}(\mathbf{x},t)$ parameterized by $\theta$ to estimate noise components, operating through sequential refinement steps denoted by a temporal parameter $t$. The denoising procedure typically involves iterative updates where the network processes corrupted inputs $\mathbf{x}$ at progressive noise levels determined by $t$.
@@ 3 Method
In this section, we delve into the specifics of our method, aiming to address the limitations of previous methods. The overall framework of our method is shown in Fig. 2.
3.1 Preliminaries
The training protocol for latent diffusion models involves two core phases. Initially, an encoder network $E(\cdot)$ compresses input samples $\mathbf{x}$ into latent representations $\mathbf{z}=E(\mathbf{x})$. These latent codes undergo systematic corruption through a diffusion trajectory determined by uniformly sampled timesteps $t$, implemented via the forward process formulation:
$\text{noise}(\mathbf{z},\epsilon,t)=\sqrt{\bar{a}{t}}\mathbf{z}+(1-\sqrt{\bar{a}{t}})\epsilon,$ (1)
where time-dependent attenuation factors $\sqrt{\bar{a}_{t}}$ govern the noise blending proportions throughout the diffusion stages.
The denoising network $\epsilon_{\theta}$ learns to recover original signals by processing corrupted latent representations paired with their corresponding timestep indices. This is achieved through the optimization objective:
$\mathcal{L}_{t}(\mathbf{z},\epsilon)=|\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),t)-\epsilon|_{2}^{2}.$ (2)
During the generation phase, novel samples are synthesized through a reverse diffusion process [29, 13] that progressively refines Gaussian noise samples $\mathbf{z}_{T}$ over multiple iterations. For conditional generation tasks like text-to-image synthesis, the framework incorporates semantic conditioning signals $c$ (e.g., text embeddings) as auxiliary inputs. This extends the denoiser architecture to $\epsilon_{\theta}(\mathbf{z},t,c)$, modifying the training objective accordingly:
$\mathcal{L}_{t}(\mathbf{z},\epsilon,c)=|\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),t,c)-\epsilon|_{2}^{2}.$ (3)
@@ 3.1 Preliminaries
0:  ($\mathcal{T}$, $\mathcal{Y}$): Real dataset and corresponding label texts.
0:  $\mathcal{P}$: Selected patches.
0:  $I$: Pre-trained image encoder.
0:  $T$: Pre-trained text encoder.
0:  $\mathcal{U}$: Pre-trained time-conditional U-Net.
0:  $N$: Number of diffusion time-steps.
0:  $C$: Number of top cluster centers.
@@
8:           $t\sim U(0.1,0.7)$
9:           $L_{i}=L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)$
10:        end for
11:        $R(\mathbf{z}|c)=\frac{1}{N}\sum_{i}L_{i}$
12:        $\mathcal{P}_{Y}\leftarrow\mathcal{P}_{Y}\cup\mathop{\arg\max}\limits_{\mathbf{p}\in\mathbf{z}}R(\mathbf{p}|c)$
13:     end for
14:     $cluster(\mathcal{U}(I(\mathcal{P}_{Y}),c,1.6))$
[…]
16:        $\mathcal{S}_{Y}\leftarrow\mathcal{S}_{Y}\cup top(\mathcal{P}_{Y}^{i})$
17:     end for
18:     $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{Y}$
19:  end for
19:  $\mathcal{S}$: Distilled dataset.
@@ 3.2 Optimal Class-relevant Feature Selection
In recent years, several methods propose the utilization of pre-trained diffusion models to directly generate images in order to improve dataset distillation performance on large-scale datasets. However, these methods often face the challenge of distributional differences between the pre-training dataset and the target dataset. Minimax [7] relies on fine-tuning a diffusion model pre-trained on ImageNet-1K, while D4[30] attempts to obtain representative samples through clustering in latent space. Both of them remain constrained by the distributional shifts introduced during the unconstrained denoising process.
To address the issue of data heterogeneity, we introduce a framework consisting of two stages that utilizing the diffusion model for localization rather than generation. As illustrated in the Stage I of Fig. 2, we first utilizes diffusion models as a zero-shot image-wise classifiers similar to recent work [18], which can be formulated as:
$p(c_{i}|\mathbf{z})$ $=\frac{\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),\;t,\;c_{i})-\epsilon]\}}{\sum_{j}\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),\;t,\;c_{j})-\epsilon]\}}$ (4)
$=\frac{1}{\sum_{j}\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c_{i})-L_{t}(\mathbf{z},\epsilon,c_{j})]\}}.$
where $c_{i}\in[C]$ and $[C]$ represents the label space, $\mathbf{z}$ is the latent of input image $\mathbf{x}$. However, such an approach can only select the image with the highest classification probability, lacking further constraints. Additionally, a larger class space can lead to significant computational overhead. Using the classifier-free guidance[12], we redefine the label space for a specific label $c$ and rewrite Eq. 4 as follow:
$p(c|\mathbf{z})=\frac{1}{1+\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)]\}},$ (5)
where $\phi$ and $c$ represents the classifier-free guidance and label text prompt respectively (i.e., ”An image of .” and ”An image of c”). To simplify computation, we define the representative of an image as follow:
$R(\mathbf{z}|c)=\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)].$ (6)
Obviously Eq. 5 and Eq. 6 are monotonically equivalent. Based on this, to further reduce the complexity of the dataset, we calculate the $R(\mathbf{p}|c)$ within each patch $\mathbf{p}$ in image and select the most typical patches as the synthetic dataset. The detailed workflow is depicted in Algorithm 1.

### v4

# Towards Consistent and Efficient Dataset Distillation via Diffusion-Driven Selection

@@ Abstract
Dataset distillation provides an effective approach to reduce memory and computational costs by optimizing a compact dataset that achieves performance comparable to the full original. However, for large-scale datasets and complex deep networks (e.g., ImageNet-1K with ResNet-101), the vast optimization space hinders distillation effectiveness, limiting practical applications. Recent methods leverage pre-trained diffusion models to directly generate informative images, thereby bypassing pixel-level optimization and achieving promising results. Nonetheless, these approaches often suffer from distribution shifts between the pre-trained diffusion prior and target datasets, as well as the need for multiple distillation steps under varying settings. To overcome these challenges, we propose a novel framework that is orthogonal to existing diffusion-based distillation techniques by utilizing the diffusion prior for patch selection rather than generation. Our method predicts noise from the diffusion model conditioned on input images and optional text prompts (with or without label information), and computes the associated loss for each image-patch pair. Based on the loss differences, we identify distinctive regions within the original images. Furthermore, we apply intra-class clustering and ranking on the selected patches to enforce diversity constraints. This streamlined pipeline enables a one-step distillation process. Extensive experiments demonstrate that our approach consistently outperforms state-of-the-art methods across various metrics and settings.
@@ 1 Introduction
Figure 1: Comparison of diffusion-based distillation methods. (a) Fine-tuning a diffusion model pre-trained on ImageNet-1K to directly generate images. (b) Fine-tuning the prototypes of the original images in latent space to generate images. (c) Identifying the most representative regions in the real images. The images generated by our method more effectively represent the feature distribution of the original dataset.
@@
The rapid advancement of deep learning has driven impressive performance gains through increasingly deeper and more complex models trained on large-scale datasets [he2016deep, dosovitskiy2020image]. However, training these models demands significant memory and computational resources. Dataset distillation has recently emerged as an effective approach to data compression [lei2023comprehensive], offering substantial reductions in resource consumption. By optimizing images in pixel space through information matching (e.g., gradient matching), these optimization-based methods produce smaller synthetic datasets that can achieve comparable performance to the original data [zhao2021dsa, zhao2021datasetdm, cazenavette2022dataset]. Despite their promising performance, the vast pixel-level optimization space restricts the feasibility of these methods to small-scale datasets, e.g., CIFAR-10/100 [krizhevsky2009learning], thus limits their generalization. A recent approach, SRe2[yin2024squeeze], scales optimization-based methods to larger datasets (e.g., ImageNet-1K [deng2009imagenet]) by using a pre-trained classifier to guide dataset synthesis. However, its strong dependence on proxy models during optimization often hinders the distilled dataset’s generalization, limiting broader applicability. Another critical limitation arises when handling various distillation configurations, such as different subsets of large datasets. Previous optimization-based methods typically incur substantial redundant costs due to repeated distillation requirements. Although recent works [he2024multisize, he2024you] aim to address this issue, their applicability remains confined to small-scale datasets.
@@
Subsequent methods [sun2024diversity, gu2024efficient, su2024d, chen2025influence, chan2025mgd, zhao2025taming] that avoid direct image optimization have shown promise for scaling dataset distillation to larger datasets and more complex model architectures, addressing several severe limitations of previous methods. RDED [sun2024diversity] leverages a pre-trained classifier (e.g., ResNet-18 [he2016deep]) to select patches randomly cropped from real images and then concatenate them into synthetic images. While RDED achieves substantial performance improvements, it requires careful model selection and retraining of a classifier tailored to the specific dataset. Moreover, concatenated images often lack complete class-relevant features, especially for complex models with higher learning capacity. Model-Finetuned methods [gu2024efficient, chen2025influence] address distillation by pruning the DiT model [peebles2023scalable] to generate synthetic datasets using regularization based on dataset representativeness and diversity. However, these methods rely on a diffusion model pre-trained on ImageNet-1K, requiring fine-tuning for different target datasets, which limits flexibility. Prototype-Finetuned methods [su2024d, chan2025mgd] apply a pre-trained text-to-image Stable Diffusion (SD)[rombach2022high] to generate synthetic datasets by finetuning original images in latent space. However, without constraints on distributional differences between the LDM’s pre-training data (e.g., LAION[schuhmann2022laion]) and the target dataset (e.g., ImageNet-1K), these methods can produce class-irrelevant or noisy images.
@@
To overcome these limitations, we propose a novel dataset distillation paradigm that leverages diffusion model’s prior to select the most informative image patches. Specifically, for each input image from the original dataset, we first predict the noise map using the pre-trained SD, with and without the label text prompt, calculating losses for each scenario. The differential loss quantifies representativeness, enabling us to crop patches most relevant to the label guided by the target dataset’s data distribution. Additionally, we can effectively constrain the implicit data distribution learned by the diffusion model by using the pre-trained SD to identify class-pertinent regions.
@@
To mitigate the inherent diversity of diffusion models, we apply clustering to capture the visual features most representative of each class. Specifically, we compute diffusion features for all patches inspired by recent work [tang2023emergent], allowing for efficient clustering and ranking of the cropped patches. Figure 1 illustrates a comparison between our method and previous diffusion-based generative distillation methods. Extensive experiments demonstrate that our approach achieves state-of-the-art (SOTA) performance on large-scale datasets, supporting an efficient, unrestricted, one-step distillation process.
@@
  • Our proposed paradigm enables a highly efficient, one-step distillation process, eliminating the need for model training and fine-tuning, and avoiding repetitive distillation across class combinations or compression ratios.
  • Extensive experiments demonstrate that our method consistently surpasses existing training-free methods across various settings.
@@ 2.1 Dataset Distillation
Dataset distillation was initially regarded as a meta-learning problem [wang2018dataset]. It involves minimizing the loss function of the synthetic dataset through a model trained on this dataset. Since then, several approaches have been proposed to enhance the performance of dataset distillation.
Recent methods propose the decoupled distillation strategy and successfully extend the dataset distillation application to large-scale datasets. SRe2[yin2024squeeze] leverages a pre-trained classifier to compute the cross-entropy loss of the synthetic dataset for updates, while RDED [sun2024diversity] calculates the confidence of randomly cropped patches from the original images with a pre-trained classifier and concatenates them together to form the synthetic dataset. Recent methods [cazenavette2023generalizing, zhong2024hierarchical] have introduced GAN as a parameterization technique. While with the development of generative models, the diffusion models are introduced to directly generate synthetic datasets, achieving promising performance [gu2024efficient, su2024d, chen2025influence, zhao2025taming, chan2025mgd]. However, existing diffusion-based methods are constrained by the implicit data distribution learned by diffusion models, leading to significant distributional bias in the synthetic dataset. In addition, aforementioned methods still fail to complete the distillation process in one pass for all settings. To address these issues, we utilize diffusion model to actively identify the most class-relevant regions in the original images, enabling an efficient dataset distillation process.
@@ 2.2 Diffusion Model
As a class of generative approaches, diffusion-based frameworks learn to progressively convert Gaussian-distributed noise vectors $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into samples matching target distributions $\mathcal{X}$, demonstrating remarkable effectiveness across diverse applications [zhu2024multibooth, zhu2024instantswap, siglidis2024diffusion]. These models employ a neural network $\epsilon_{\theta}(\mathbf{x},t)$ parameterized by $\theta$ to estimate noise components, operating through sequential refinement steps denoted by a temporal parameter $t$. The denoising procedure typically involves iterative updates where the network processes corrupted inputs $\mathbf{x}$ at progressive noise levels determined by different timestap $t$.
The training protocol for latent diffusion models involves two core phases. Initially, an encoder network $E(\cdot)$ compresses input samples $\mathbf{x}$ into latent representations $\mathbf{z}=E(\mathbf{x})$. These latent codes undergo systematic corruption through a diffusion trajectory determined by uniformly sampled timestep $t$, implemented via the forward process formulation:
$N_{t}(\mathbf{z},\epsilon,t)=\sqrt{\bar{a}{t}}\mathbf{z}+(1-\sqrt{\bar{a}{t}})\epsilon,$ (1)
where $N_{t}$ denotes the predicted noise at timesteps $t$, and the time-dependent attenuation factors $\sqrt{\bar{a}_{t}}$ govern the noise blending proportions. The denoising network $\epsilon_{\theta}$ then learns to recover original signals, which is achieved through the optimization objective:
$\mathcal{L}_{t}(\mathbf{z},\epsilon)=|\epsilon_{\theta}(N_{t}(\mathbf{z},\epsilon,t),t)-\epsilon|_{2}^{2}.$ (2)
During the generation phase, novel samples are synthesized through a reverse diffusion process [song2020denoising, ho2020denoising] that progressively refines Gaussian noise samples $\mathbf{z}_{T}$ over multiple iterations. For conditional generation tasks like text-to-image synthesis, the framework incorporates semantic conditioning signals $c$ (e.g., text embeddings) as auxiliary inputs. This extends the denoiser architecture to $\epsilon_{\theta}(\mathbf{z},t,c)$, modifying the training objective accordingly:
$\mathcal{L}_{t}(\mathbf{z},\epsilon,c)=|\epsilon_{\theta}(N_{t}(\mathbf{z},\epsilon,t),t,c)-\epsilon|_{2}^{2}.$ (3)
@@ 3 Method
In this section, we delve into the specifics of our method, aiming to address the limitations of previous methods. The overall framework of our method is shown in Figure 2.
@@
0: ($\mathcal{T}$, $\mathcal{Y}$): Real dataset and corresponding label texts. $\mathcal{P}$: Selected patches. II: Pre-trained image encoder. $T$: Pre-trained text encoder. $\mathcal{U}$: Pre-trained time-conditional U-Net. $T$: Number of diffusion time-steps. $C$: Number of top cluster centers.
@@
7:   for i $\leftarrow$} 0 to N1N-1 do
8:    $t\sim U(0.1,0.7)$
9:    $N_{i}=N_{t}(\mathbf{z},\epsilon,c)-N_{t}(\mathbf{z},\epsilon,\phi)$
10:   end for
11:   $M(\mathbf{z}|c)=\frac{1}{T}\sum_{i}N_{i}$
12:   $\mathbf{p}=\mathop{\arg\max}\limits_{\mathbf{p}}\sum_{\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbf{p}}\hat{M}(\mathbf{x}|c)[i,j])$
13:   $\mathcal{P}_{Y}\leftarrow\mathcal{P}_{Y}\cup\{\mathbf{p}\}$
14:  end for
15:  $cluster(\mathcal{U}(I(\mathcal{P}_{Y}),c,0.16))$
16:  for i $\leftarrow$} 0 to C1C-1 do
17:   $\mathcal{S}_{Y}\leftarrow\mathcal{S}_{Y}\cup top(\mathcal{P}_{Y}^{i})$
18:  end for
19:  $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{Y}$
20:end for
@@ 3.1 Optimal Class-relevant Feature Selection
In recent years, several methods propose the utilization of pre-trained diffusion models to directly generate images in order to improve dataset distillation performance on large-scale datasets. However, these methods often face the challenge of distributional differences between the pre-training dataset and the target dataset. Model-finetuned methods [gu2024efficient, chen2025influence] rely on fine-tuning a diffusion model pre-trained on ImageNet-1K, while Prototypes-finetuned methods [su2024d, chan2025mgd] attempt to obtain representative samples through finetuning the original images in latent space. Both of them remain constrained by the distributional shifts introduced during the unconstrained denoising process.
To address the issue of data heterogeneity, we introduce a framework consisting of two stages that utilizing the diffusion model for localization rather than generation. As illustrated in Stage I of Figure 2, we first utilize diffusion models as zero-shot image-wise classifiers [li2023your], which can be formulated as:
$p(c_{i}|\mathbf{z})$ $=\frac{\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(N_{t}(\mathbf{z},\epsilon,t),\;t,\;c_{i})-\epsilon]\}}{\sum_{j}\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(N_{t}(\mathbf{z},\epsilon,t),\;t,\;c_{j})-\epsilon]\}}$ (4)
$=\frac{1}{\sum_{j}\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[N_{t}(\mathbf{z},\epsilon,c_{i})-N_{t}(\mathbf{z},\epsilon,c_{j})]\}}.$
where $c_{i}\in[C]$ and $[C]$ represents the label space, $\mathbf{z}$ is the latent of input image $\mathbf{x}$. However, such an approach can only select the image with the highest classification probability, lacking further constraints. Additionally, a larger class space can lead to significant computational overhead. Using the classifier-free guidance[ho2022classifier], we redefine the label space for a specific label $c$ and rewrite Equation 4 as follow:
$p(c|\mathbf{z})=\frac{1}{1+\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[N_{t}(\mathbf{z},\epsilon,c)-N_{t}(\mathbf{z},\epsilon,\phi)]\}},$ (5)
where $\phi$ and $c$ represents the classifier-free guidance and label text prompt respectively (i.e., “An image of .” and “An image of c”). We then simplify the computation as follow:
$M(\mathbf{z}|c)=\mathbb{E}_{\epsilon,t}[N_{t}(\mathbf{z},\epsilon,c)-N_{t}(\mathbf{z},\epsilon,\phi)].$ (6)
Obviously Equation 5 and Equation 6 are monotonically equivalent. Based on the differential noise map $M(\mathbf{z}|c)\in\mathbb{R}^{H\times W\times 3}$, we could calculate the representative score $R(\mathbf{p}|c)$ within each patch $\mathbf{p}$ in original image. To enable computation on the original image, we first upsample the noise map from the latent space to the original image resolution via interpolation and take the channel-wise average to obtain $\hat{M}(\mathbf{x}|c)\in\mathbb{R}^{H\times W}$. We then apply an average pooling kernel to compute the score for each patch, as illustrated below:
$R(\mathbf{p}|c)=\sum_{\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbf{p}}\hat{M}(\mathbf{x}|c)[i,j].$ (7)
Based on these representation scores, we can select the final distilled data and further reduce the complexity. The detailed workflow is depicted in Algorithm 1.